zarkam21:

What is the volume of a piece of material where a 77.3g sample has a density of 0.630 g/cm^3

Vocaloid:

density = mass / volume so plugging density and mass in 0.0630 = 77.3/x solve for volume

zarkam21:

plug in 0.0630 or 0.630?

Vocaloid:

oh 0.630 not 0.0630

Vocaloid:

0.630 = 77.3/x

kittybasil:

It's 0.630, that was a typo \[D=\frac{m}{v}\]D = \(0.630\text{ g/cm}^3\) m = \(77.3\text{ g}\)\[\therefore0.630\text{ g/cm}^3=\frac{77.3\text{ g}}{v}\]\[v=77.3\text{ g}\div0.630\text{ g/cm}^3=??\]

zarkam21:

122

Vocaloid:

almost, it ends up being 122.698 which rounds up to 123 cm^3 not 122

kittybasil:

I was about to say `don't forget units` (but yes what Voca said)

zarkam21:

wait units are L

Vocaloid:

hm? the units they gave you for density are 0.630 g/cm^3 so the volume itself is cm^3

kittybasil:

I mean... density can also be `g/mL` but it depends What units were you given?\[\text{What is the volume of a piece of material where a }77.3g\]\[ \text{sample has a density of }0.630 g/cm^3\]I don't see liters mentioned here

kittybasil:

But if you need to convert... Try this